OpenStudy (anonymous):

what is the probability of pulling at least 1 ace from a 52 card deck

OpenStudy (perl):

what do you mean by 'at least'. are you drawing 5 cards?

OpenStudy (anonymous):

oh sorry im pulling 3

OpenStudy (perl):

ok. We can solve this using the complement method. P ( at least one ace in three draws ) = 1 - P( no ace in three draws )

OpenStudy (perl):

how many 'non ace' cards are there?

OpenStudy (perl):

$$ \Large { P(A') = 1 - P(A) \\ \therefore \\ P(A) = 1 - P(A~' ) } $$

OpenStudy (anonymous):

thanks!!